The Gourmand Cooking School runs short cooking courses at its small campus. Management has identified two cost drivers that it uses in its budgeting and performance reports—the number of courses and the total number of students. For example, the school might run two courses in a month and have a total of 61 students enrolled in those two courses. Data concerning the company’s cost formulas appear below: Fixed Cost per Month Cost per Course Cost per Student Instructor wages $ 2,940 Classroom supplies $ 270 Utilities $ 1,200 $ 50 Campus rent $ 4,700 Insurance $ 2,400 Administrative expenses $ 3,800 $ 44 $ 3 For example, administrative expenses should be $3,800 per month plus $44 per course plus $3 per student. The company’s sales should average $900 per student. The actual operating results for September appear below: Actual Revenue $ 52,000 Instructor wages $ 11,040 Classroom supplies $ 16,320 Utilities $ 1,810 Campus rent $ 4,700 Insurance $ 2,540 Administrative expenses $ 3,585 Required: 1. The Gourmand Cooking School expects to run four courses with a total of 61 students in September. Complete the company’s planning budget for this level of activity. 2. The school actually ran four courses with a total of 51 students in September. Complete the company’s flexible budget for this level of activity. 3. Calculate the revenue and spending variances for September. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1
1
Planning Budget
For the month ended September 30
Revenue 54900 =61*900
Expenses:
Instructor wages 11760 =4*2940
Classroom supplies 16470 =61*270
Utilities 1400 =1200+(4*50)
Campus rent 4700
Insurance 2400
Administrative expenses 4159 =3800+(4*44)+(61*3)
Total expense 40889
Net operating income 14011
2
Flexible Budget
For the month ended September 30
Revenue 45900 =51*900
Expenses:
Instructor wages 11760 =4*2940
Classroom supplies 13770 =51*270
Utilities 1400 =1200+(4*50)
Campus rent 4700
Insurance 2400
Administrative expenses 4129 =3800+(4*44)+(51*3)
Total expense 38159
Net operating income 7741
3
Revenue and spending variance
Revenue 6100 F
Expenses:
Instructor wages 720 F
Classroom supplies 2550 U
Utilities 410 U
Campus rent 0 None
Insurance 140 U
Administrative expenses 544 F
Total expense 1836 U
Net operating income 4264 F
Workings for 3:
Actual results Revenue and spending variance Flexible budget
Courses 4 4
Students 51 51
Revenue 52000 6100 F 45900
Expenses:
Instructor wages 11040 720 F 11760
Classroom supplies 16320 2550 U 13770
Utilities 1810 410 U 1400
Campus rent 4700 0 None 4700
Insurance 2540 140 U 2400
Administrative expenses 3585 544 F 4129
Total expense 39995 1836 U 38159
Net operating income 12005 4264 F 7741